OpenStudy (anonymous):

someone pleaseeee help, .....jennifer invested $4,000 in her savings account for 4 years. when she withdrew it, she had 4,731.75. interest is compounded continuously. what was the interest rate on the account?

OpenStudy (amistre64):

A = Pe^rt A/P = e^rt ln(A/P) = rt ln(A/P) ------ = r t

OpenStudy (anonymous):

im not understanding where to plug inthe numbers

OpenStudy (amistre64):

when we take a principal amount, and compound it ... it grows bigger; that should help you define A and P

OpenStudy (amistre64):

t is just the number of years

OpenStudy (anonymous):

i got 4.2% is that what you got?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

thank u soo much